40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 26 01:00:23 2017, gautam, Update, General, DRMI locking  DRMI_1f_26April2017.pngIMG_7373.JPGIMG_7372.JPG
    Reply  Fri Apr 28 02:04:36 2017, gautam, Update, General, DRMI locking  20170428_DRMI.pdf
       Reply  Fri Apr 28 13:56:26 2017, rana, Update, General, DRMI locking  
          Reply  Mon May 1 16:29:51 2017, gautam, Update, General, DRMI locking  LSC_sensingNoise.pdfIMG_7378.JPG
             Reply  Mon May 1 21:45:54 2017, ericq, Update, General, DRMI locking  
                Reply  Wed May 3 18:45:45 2017, rana, Update, General, DRMI locking  
                   Reply  Thu May 4 19:03:15 2017, gautam, Update, General, DRMI locking - preliminary MICH NB DRMI_noArms_April30.pdfMICH_OLTF.pdfC1NB_disp_40m_MICH_NB_30_April_2017.pdf40m_REFL_RFPDs_efficiency.pdf
                      Reply  Fri May 5 10:13:02 2017, ericq, Update, General, MICH NB questions 
                         Reply  Fri May 5 12:10:53 2017, gautam, Update, General, MICH NB questions C1NB_disp_40m_MICH_NB_30_April_2017.pdf
                            Reply  Sat May 6 21:52:11 2017, rana, Update, General, MICH NB questions 
                               Reply  Wed May 10 01:56:06 2017, gautam, Update, General, MICH NB - OL coupling OL_BS_ITM_comp.pdfC1NB_disp_40m_MICH_NB_8_May_2017.pdfC1NB_disp_40m_MICH_NB_10_May_2017.pdf
                                  Reply  Wed May 10 16:53:38 2017, rana, Update, General, MICH NB - OL coupling 
                                  Reply  Wed May 10 17:17:05 2017, gautam, Update, General, DAC / Coil Driver noise actuation.jpg
                                     Reply  Wed May 10 17:46:44 2017, gautam, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                        Reply  Thu May 11 09:45:46 2017, rana, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                           Reply  Thu May 11 18:59:22 2017, gautam, Update, General, SRM coil driver + dewhite board initial survey D010001-B_40m.pdfD000183-C8_40m.pdf
                                              Reply  Fri May 12 01:36:04 2017, gautam, Update, General, SRM coil driver + dewhite board LISO modeling SRM_bypass_plus_CoilDriver.pdfliso.zip
                                                 Reply  Fri May 12 12:34:55 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                    Reply  Fri May 12 18:50:08 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                       Reply  Mon May 22 13:37:01 2017, gautam, Update, General, DAC noise estimate DAC_noise_model.pdfC1NB_disp_40m_MICH_NB_22_May_2017.pdf
                                                          Reply  Tue May 23 22:58:23 2017, gautam, Update, General, De-Whitening board noises ITMX_deWhite_ch1_noise.pdf
                                                             Reply  Thu May 25 19:27:29 2017, gautam, Update, General, Coil driver board noises  coilDriverNoises.pdf
                                                                Reply  Mon May 29 16:47:38 2017, gautam, Update, General, Coil driver boards reinstalled deWhitening_consolidated.pdfdeWhitening_measurements.zipETMY_OL.png
                                                                   Reply  Tue May 30 16:02:59 2017, gautam, Update, General, Coil driver boards reinstalled OL_noiseInjection.pdf
                                                                      Reply  Thu Jun 1 00:10:15 2017, gautam, Update, General, Coil driver boards reinstalled 
       Reply  Fri Apr 28 19:32:06 2017, gautam, Update, General, DRMI locking - PRCL angular FF PRC_angularFF.pdfPRC_TFs.pdf
Message ID: 12951     Entry time: Wed Apr 26 01:00:23 2017     Reply to this: 12954
Author: gautam 
Type: Update 
Category: General 
Subject: DRMI locking  

Since we'd like to get back to DRSE locking, I tried locking the DRMI tonight. I did the following:

  • First, I aligned the arms, and ran the dither alignment scripts to maximize the arm transmission
  • Next, I misaligned the ETMs, and tried to lock the PRC resonant for the carrier (i.e. PRCL on REFL11I, MICH on AS55Q). I got brief lock stretches of a few seconds but not longer. Turns out the AS55 beam was barely hitting the photodiode. I guess this wasn't looked at since Johannes modified the AS path for the loss measurements. Anyways, it just required a minor tweak to center the beam on the AS55 photodiode.
  • Once the PRC was locked, I ran the PRC and MICH dither align scripts. The way these are set up right now, the error signals to these servos are REFLDC and ASDC respectively (demodulated at the respective dither frequencies). But looking at the spots on the ITM cameras with the PRC resonant, the spots seem shifted (in both PIT and YAW) relative to the spots when the arm cavity is resonant. Shouldn't they be the same mode? Or maybe I am missing something.
       
  • Next, I tried to lock the DRMI with the 1f error signals: i.e. PRCL on REFL11 I, SRCL on REFL55 I, and MICH on AS55 Q. After some demod phase tweaking, I was able to get some locks going. Turning on the PRC angular feedforward seemed to help the locking, but I have no idea if the installed filters are still the correct ones. I believe the POP QPD channels are the witnesses used to train this filter, I will look at the predicted vs achieved subtraction.
  • At this point, I was able to get locks lasting a few minutes - see the attachment. I ran the UGF servos and tweaked the loop gains a little, but before I could start a loop measurement, I lost the lock. I am calling it for the night.

GV 26 April 2017, 3pm: Forgot to note yesterday that I re-connected the suspect Satellite box, which has been connected to the SRM signal chain, back to the SRM suspension. I did not see any instances of glitching during my work last night. Also added pictures showing shifted spots on ITMs when PRC is locked relative to when arms are locked...

ELOG V3.1.3-